# Decoupled Data Consistency with Diffusion Purification for Image Restoration

### v5

@@ 5.2.1 Pixel Space Diffusion Models
5.2.2 Latent Space Diffusion Models
For the results with LDMs, we display quantitative results in Table 1 and Table 2 highlighted in blue, with qualitative results in Figure 7. Recall that as outlined in Section 4.2, there are two approaches for ensuring data consistency. For the best performance, we use pixel space optimization (Approach 2) for motion deblurring and Gaussian deblurring and latent space optimization (Approach 1) for box inpainting and super-resolution. As observed in the tables, we observe that our algorithm outperforms the existing baselines on many of the inverse problems. Interestingly, we observe that PSLD seems to reconstruct images that are often noisy and blurry, where the learning rate plays a huge role on the reconstruction quality. While ReSample may have higher LPIPS scores on some inverse problem tasks, our algorithm is able to achieve better PSNR and SSIM scores.
@@ 5.3 Ablation Studies
Finally, we conduct ablation studies demonstrating computational efficiency, convergence, and robustness to measurement noise.
Computational Efficiency.
@@
First, we compare the inference times (the time it takes to reconstruct a single image) of our method with the baselines and display the results in Table 3. Clearly, for pixel based diffusion models, our methods are much more efficient than DPS [16] and MCG [13], as we do not require backpropogation through the score function for data consistency. Furthermore, using the single step Tweedie’s formula approximation for diffusion purification is much faster than using reverse SDE which involves multiple time steps.
Second, for latent space methods, our methods are also faster than existing methods PSLD [20] and ReSample [21] as observed in Table 3. For our own methods, we observe that optimizing at the pixel level (Approach 2) for data consistency proves to be considerably more efficient than optimizing in the latent space (Approach 1). As previously mentioned, this is due to the need for backpropagating through the complex deep decoder $\mathcal{D}(\cdot)$. This is also one of the main reasons our methods are more computationally efficient than methods such as ReSample.
Algorithm Time (sec)
DPS [16] 387
MCG [13] 385
Our Pixel (Reverse SDE, Version 1) 28.7
Our Pixel (Tweedie’s Formula, Version 2) 2.14
PSLD [20] 188
ReSample [21] 118
Our Latent (Latent DC + Reverse SDE) 109
Our Latent (Pixel DC + Reverse SDE) 12.2
Our Latent (Latent DC + Tweedie’s Formula) 100
Our Latent (Pixel DC + Tweedie’s Formula) 2.59
Table 3: Comparison of inference times. The results colored in blue are for LDMs, whereas the others are for pixel space diffusion models.
@@
Figure 8: Convergence of our methods in terms of LPIPS. The orange, green, red and blue curves correspond to the reconstruction process with data fidelity optimization only, our proposed method with a constant (large) purification strength, constant (small) purification strength and a linear decaying purificaiton strength, respectively. The x-axis denotes the total number of data fidelity gradient steps.
Convergence Study.
@@
In this section, we investigate the convergence of our algorithm in terms of LPIPS. This study also serves to demonstrate the efficiency of a linear decaying time schedule for $T_{k}$ for diffusion purification across various image restoration problems. As shown in Figure 8, at the beginning of each diffusion purification and data fidelity optimization cycle, the LPIPS increases after performing diffusion purification due to stochasticity introduced by the forward-backward diffusion process. However, the subsequent iterations can successfully decrease LPIPS and result in reconstructions with higher qualities compared to pure data fidelity optimization, illustrating that the resulting purified image is able to serve as a good initialization point. In the same figure, also notice that the linear decaying schedule greatly enhances the performance of our algorithm compared to using a constant time schedule for $T_{k}$. This is expected, as a large constant purification strength will result in serious hallucination while a small constant purification strength is not able to fully alleviate the artifacts in the data fidelity based estimations. In Figure 9, we illustrate an example reconstruction process for super-resolution with noisy measurements, with more examples in the Appendix. It is evident that the quality of the reconstructed image gradually increases with more iterations.
@@
Metric DPS [16] Version 1 Version 2 DPS [16] Version 1 Version 2
@@
Thus far, we considered experiments on image restoration tasks in which the measurements were noiseless. However, in many practical settings of interest, the measurements $\bm{y}$ can be corrupted with additive noise. In this section, we perform an ablation study on the performance of our algorithm in which the measurements have additive Gaussian noise $\mathcal{N}(0,\sigma^{2}_{\bm{y}})$, with varying values of $\sigma_{\bm{y}}$ for super resolution and motion deblurring. We compare our algorithm to DPS, as their theory states that the approximation of their gradient likelihood term actually becomes tighter as the measurement noise $\sigma_{\bm{y}}$ increases [16]. We conduct experiments on $100$ randomly selected images from the LSUN-Bedroom dataset for super-resolution, motion deblurring on standard deviations $0.05,0.075$, and $0.1$. For the hyperparameters, we tune the parameters according to $\sigma_{\bm{y}}=0.01$ and fix them throughout the experiments. The rest of the setup is the same as previously described. We present the results in Table 4. Here we observe that as the standard deviation increases, the performance of all algorithms slightly degrades, as expected. However, our algorithm consistently outperforms DPS in metrics such as PSNR and SSIM, whereas DPS often exhibits higher perceptual quality measured by LPIPS. This difference arises because DPS frequently suffers from hallucinations, introducing artifacts in the image that “sharpen” it. Our algorithm is less susceptible to such artifacts, resulting in “smoother” images with higher signal-to-noise ratios. Details of the hyperparameters and visual results can be found in Section A.3.
6 Relationship to Prior Arts
Here, we discuss the relationship of our work with existing methods for image restoration, including both methods derived in the pixel space and algorithms tailored for the latent space.
Pixel Space Diffusion Models.
Other than DPS [16], there are other diffusion model-based inverse problem solvers available for addressing image restoration tasks. Song et al. [17] introduced $\Pi$GDM, an algorithm that approximates the likelihood gradient using a pseudo-inverse of the linear sensing matrix used for generating the measurements. When the operator $\mathcal{A}(\cdot)$ is linear, it can be represented by a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, for which one can calculate its pseudo-inverse. They leverage this property by first constructing the conditional distribution $p(\bm{y}|\bm{x}_{t})$ in terms of the sensing matrix $\bm{A}$ and derive the likelihood gradient using this conditional distribution. The work by Wang et al. [15] has a similar flavor; since the image restoration problem is ill-posed, the reconstructions without priors generally are a mixture of a sample in both the range and the nullspace of the matrix $\bm{A}$. They use the diffusion model as a prior to essentially “fill in” this nullspace, which is constructed by taking the pseudo-inverse of $\bm{A}$ and used during the reverse sampling process. The work presented in [50] introduced the Direct Diffusion Bridge method building upon the DPS approach. Here, the data fidelity is enforced during the reverse sampling procedure by taking a deeper gradient step which is designed for improved data consistency.
Another popular method for addressing image restoration problems is the Denoising Diffusion Restoration Models (DDRMs) proposed by Kawar et al. [51]. DDRMs construct a probabilistic graph conditioned on the measurements, allowing the use of a pre-trained unconditional diffusion model. Subsequently, they perform the reverse sampling process in the “spectral space”, which is constructed by taking the SVD of the sensing matrix $\bm{A}$. These works generally share two common themes: they require the inverse problem to be linear, and they incorporate the data consistency steps in the reverse diffusion process. Although we do not test on nonlinear inverse problems, our algorithm does not face these limitations and is flexible in this regard.
Lastly, the study presented by Luther et al. [52] shares close ties with Version 2 our algorithm in the pixel space. They introduced Diffusive Denoising of Gradient-based Minimization (DDGM), an image-space method for solving linear tomographic reconstruction problems, which iteratively applies data consistency steps and image denoising through a pre-trained diffusion model. However, as we have shown in Section 5.2, directly regularizing images with Tweedie’s formula will result in blurry images. In contrast, the proposed Version 1 of our algorithm alleviates this issue by performing diffusion purification with multiple diffusion reverse sampling steps. Furthermore, we propose a unified framework that can incorporate accelerated samplers, latent diffusion models, and consistency models, thereby significantly reducing the inference time. Hence, our work addresses a broader spectrum, not only limited to tomographic reconstruction, and can be readily adapted to applications requiring fast inference.
Latent Space Diffusion Models.
Throughout our experiments, we also presented results using LDMs, using PSLD [20] and ReSample [21] as baselines. ReSample [21] is an algorithm designed to solve inverse problems in the latent space that involves solving a data fidelity optimization problem during the reverse sampling process, a process they term “hard data consistency”. The data fidelity objective is similar to the approaches we take, where they alternate between minimizing the objective in either the latent space or the pixel space. Upon solving this objective, they devise a resampling scheme to project the data-consistent sample back onto the noisy manifold to continue the reverse sampling process. Nonetheless, by decoupling these two processes, our algorithm does not require a resampling stage. Additionally, ReSample cannot use accelerated samplers because of the coupling, as it would hinder the data fidelity stage by reducing the number of sampling steps, unlike our framework.
Posterior Sampling with Latent Diffusion (PSLD) [20] extends the idea of DPS to the latent space by addressing the challenge that the decoder and encoder constitute a many-to-one mapping. This implies that there may be multiple latent codes that fit the measurements, and they devise a gradient scheme to ensure the “correct” latent code is estimated during the reverse sampling process. However, their technique involves two gradient computations—one through the LDM decoder and one through the pre-trained LDM score function. Thus, PSLD suffers from increased inference time. Furthermore, it is observed in our experiment that the performance of PSLD is sensitive to the choice of hyperparameters.
Plug-and-Play (PnP) Methods.
The Plug-and-Play Priors (PnP) framework combines physical and learned models for effective computational imaging. PnP alternates between minimizing a data-fidelity term for enforcing data consistency and applying a learned regularizer, typically an image denoiser [53, 36, 37]. The PnP framework is commonly formulated using the Alternating direction method of multipliers (ADMM) [36] and fast iterative shrinkage/thresholding algorithms (FISTA) [54]. Our proposed method shares similarity with the PnP in the sense that both methods utilize image denoisers as the implicit prior. However, while PnP performs denoising directly on $\bm{x}_{k}$ by assuming that $\bm{x}_{k}$ contains certain pseudo noise, our proposed method uses diffusion models to explicitly add noise to $\bm{x}_{k}$ and then denoises it by either iterative sampling or Tweedie’s formula [52] estimation.
7 Discussion & Conclusion
Throughout this paper, we have discussed both the efficiency and flexibility of our algorithm, demonstrating that we can compete with many state-of-the-art techniques while offering numerous advantages. However, despite these strengths, our algorithm has a few steps that warrant further investigation. Recall that the input to the diffusion purification stage is an initial reconstruction estimated from minimizing the data fidelity objective. Notice that if this step does not return an estimate that is meaningful, then the diffusion purification stage cannot produce a high-quality reconstruction. Although this was not an issue for the image restoration problem that we considered, it may be the case for more complex problems, as simply minimizing the data fidelity objective in Equation (10) can lead to poor local minima. For these tasks, it would be of great interest to explore replacing Equation (10) with more advanced solvers.
Furthermore, while our algorithm generally outperforms others in metrics such as PSNR and SSIM, it often only shows marginal improvements in LPIPS, which measures perceptual quality. This outcome is expected, as coupled methods like DPS prioritize the prior step more meticulously, potentially leading to hallucinations. On the other hand, by decoupling data fidelity from the prior, our algorithm places greater emphasis on data fidelity. Therefore, exploring methods to carefully balance between the prior and data consistency with decoupling is an important avenue for future research.
8 Data Availability Statement
@@
Acknowledgement
QQ, XL, SMK acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, MICDE Catalyst Grant, and a gift grant from KLA. IA and SR acknowlege support from NSF CCF-2212065.
Appendix A Implementation Details
Notation Definition
$K\in\mathbb{N}$ Number of iterations for overall algorithm
$\tau\in\mathbb{N}$ Total number of iterations for data consistency optimization
$\alpha\in\mathbb{R}$ Learning rate for data consistency optimization
$T_{k}\in\mathbb{R}$ Diffusion purification strength at the $k^{\text{th}}$ iteration
$\bm{v}_{k}\in\mathbb{R}^{n}$ Prior consistent sample at $k^{\text{th}}$ iteration
$\bm{x}_{k}\in\mathbb{R}^{n}$ Data consistent sample at $k^{\text{th}}$ iteration
Table 5: Summary of the notation used throughout this work.
A.1 Discussion on Hyperparameters
@@
Here, we discuss the hyperparameters associated with our algorithm and some typical choices we made throughout our experiments.
@@
  • Total Number of Iterations. We run the algorithm for a total of $K$ iterations, which refers to reconstructing the image by alternately applying data fidelity optimization and diffusion purification for $K$ rounds. Empirically, we observe a $K$ ranging from 10 to 20 is sufficient for generating high quality reconstructions. We suggest users start with $K=10$ and then gradually decrease it if a faster computation speed is required.
@@
  • Optimizer for Data Fidelity Optimization. Throughout our experiments, we use momentum-based gradient descent with the momentum hyperparameter set to 0.9. The optimal learning rate (LR) for data fidelity optimization varies from task to task. In general, the LR should neither be too large, which will not result in images with meaningful structures, nor too small, in which case the convergence speed of the overall algorithm will be too slow. Throughout the experiment we fix the number of gradient steps for each round of the data fidelity optimization. When the measurements are corrupted by noise, one can further perform early stopping to prevent overfitting to the noise.
@@
  • Decaying Time Schedule for Diffusion Purification. A decaying time schedule plays the key role in the success of our proposed algorithm. As described in  Section 3.3, given a data fidelity based reconstruction $\bm{x}_{k}$, the goal of diffusion purification is to generate $\bm{v}_{k}$ that shares the overall structural and semantic information of $\bm{x}_{k}$, such that it is consistent with the measurements while at the same time being free from the artifacts visible in $\bm{x}_{k}$. Unfortunately, the $\bm{v}_{k}$ resulting from diffusion purification is sensitive to the choice of $T$, i.e., the amount of noise added to $\bm{x}_{k}$. When $T$ is small, the resulting $\bm{v}_{k}$ will exhibit similar structure to $\bm{x}_{k}$ but with artifacts, as there is not enough noise to submerge these artifacts through diffusion purification. On the other hand, when $T$ is too large, though $\bm{v}_{k}$ will be free from artifacts, its semantic information and overall structure will significantly deviate from $\bm{x}_{k}$. As a result, the resulting reconstructions will not be measurement-consistent.
@@
    To this end, previous works that utilize diffusion purification [31, 14] depend on fine-tuning an optimal $T^{\star}$. However, we find that the fine-tuning strategy can be infeasible since the optimal $T^{\star}$ varies from task to task and even from image to image. To alleviate the need to identify the optimal $T^{\star}$, we implement a linear decaying schedule $\{T_{k}\}_{k\in[K]}$ for each iteration $k$. In the early iterations when $T_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though inconsistent with the measurements, will be free from artifacts and can serve as good starting point for the subsequent iteration of data fidelity optimization. As $T_{k}$ keeps decreasing, the algorithm will converge at an equilibrium point which is both consistent with the measurements and free from the artifacts. When the measurements are noiseless, we observe that for a diffusion model trained with 1000 total timesteps, a linear decaying time schedule starting from $T_{1}=400$ and ending at $T_{K}=0$ results in reconstructions with high quality for Super-Resolution, Gaussian Deblurring and Motion Deblurring. For Box Inpainting, we choose a higher initial timestep $T_{1}=700$ and $T_{1}=500$ for pixel space diffusion models and LDMs respectively to compensate the severe artifacts in the large box region. When the measurements are corrupted by noise, one may instead end at $T_{K}=T$, where $T>0$ to compensate the negative effect of the measurement noise. We demonstrate the reconstruction process for various tasks in Figures 10, 11, 12, 13 and 16. Notice that the diffusion purification removes the artifacts from the initial reconstructions obtained via data fidelity optimization especially in the early iterations and results in high quality reconstructions consistent with the measurements as timestep decreases.
@@
  • DDIM Steps for Diffusion Purification. With the help of accelerated diffusion samplers such as DDIMs [11], we can perform diffusion purification with as low as 20 unconditional sampling steps. As mentioned before, the total number of iterations $K$ ranging from 10-20 is sufficient for generating high quality reconstructions. This amounts to 200 to 400 unconditional reverse sampling steps in total, a significantly lower figure compared to DPS [16], which requires 1000 unconditional reverse sampling steps. Interestingly, we find that increasing the DDIM sampling steps does not necessarily improve the performance of diffusion purification. Therefore, we suggest users use 20 steps for diffusion purification.
  • Pixel Space Optimization and Latent Space Optimization. As discussed in Section 4.2, in the case of LDMs, one can use either Pixel Space Optimization Equation (20) or Latent Space Optimization (18) to enforce the data consistency. Depending on the specific task, one approach may generate $\bm{x}_{k}$’s with higher quality than the other. In practice, as shown in Figures 14 and 15, we observe that for motion deblurring and Gaussian deblurring, the initial reconstructions generated with Pixel Space Optimization have significantly higher quality compared to the ones generated with Latent Space Optimization. Furthermore, Pixel Space Optimization is much faster than the Latent Space Optimization since the gradient computation of the former approach does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$. To achieve the best reconstruction quality, we suggest users try both options and choose the one which gives the best performance. On the other hand, for applications that require fast reconstruction speed, we suggest using Pixel Space Optimization.
Figure 14: Comparison between Pixel Space Optimization and Latent Space Optimization for Motion Deblurring. Initial Reconstructions obtained via Pixel Space Pptimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.

### v6

@@ 4.2 Experimental Results
Comparison of Pixel Space Diffusion Models. We first test DCDP and the baselines based on pixel space diffusion models. The qualitative results are reported in Figure 3, and the quantitative results are presented in Tables 1, 2 and 3. In these tables, it is evident that our methods consistently outperform or achieve performance on par with the baselines. We note that Red-Diff [45] performs extremely well on nonlinear deblurring but is less competitive on other tasks.
When comparing DCDP-DDIM and DCDP-Tweedie, we observe that DCDP-Tweedie generally achieves higher PSNR and SSIM, which mainly measure the data fidelity. However, its reconstructions are often blurrier and contain less perceptual detail. As shown in Figure 4, DCDP-DDIM qualitatively preserves richer high-frequency content—evident in the Canny-edge visualizations—and leads to a better LPIPS score, which more aligns with the perceptual quality. This phenomenon is commonly referred to as the “perception-distortion trade-off” [51].
For the results with CMs, we present the quantitative results in Tables 1 and 2 highlighted in orange, and qualitative results in Figure 5, which demonstrate that DCDP with CMs leads to decent reconstruction quality for the inverse problems considered. This result highlights the flexibility of our method to accommodate a broad range of diffusion-based models. In contrast, none of the baselines considered are able to utilize CMs, as they require a sufficiently large number of reverse sampling steps. Lastly, we note that the CMs used in this work are adopted from  [17], which is the first ever prototype that introduces the concept of CMs. The performance of DCDP with CMs could potentially be further improved by incorporating more advanced consistency models proposed in recent works such as sCMs [52], Latent Consistency Models [53] and InstaFlow [54], which we leave for future investigation.
@@
Comparison of Latent Space Diffusion Models. For the results with LDMs, we present quantitative results in Tables 1, 2 and 3 highlighted in blue, and qualitative results in Figure 6. Recall that, as outlined in Section 3.4, there are two approaches for enforcing data consistency. We use pixel space data consistency (Approach II) for motion deblurring and Gaussian deblurring, and latent space optimization (Approach I) for box inpainting, super-resolution, and nonlinear deblurring, as doing so empirically leads to the best reconstruction quality. As shown in the tables, our algorithm consistently matches or outperforms all existing baselines. Interestingly, we observe that PSLD seems to reconstruct images that are often noisy and blurry, where the learning rate plays a huge role in the reconstruction quality. While ReSample could achieve higher LPIPS scores on some inverse problem tasks, our algorithm is able to achieve better PSNR and SSIM scores.
Method Nonlinear Deblurring
[…]
DPS [14] $0.355\pm 0.09$ $22.60\pm 3.36$ $0.588\pm 0.16$
RED-Diff [45] $\textbf{0.211}\pm 0.05$ $\textbf{29.51}\pm 0.76$ $\textbf{0.828}\pm 0.08$
DCDP-DDIM (Ours) $0.323\pm 0.06$ $24.08\pm 2.42$ $0.634\pm 0.112$
DCDP-Tweedie (Ours) $\underline{0.320}\pm 0.08$ $\underline{25.36}\pm 2.77$ $\underline{0.712}\pm 0.10$
Latent-DPS $0.483\pm 0.08$ $21.77\pm 3.02$ $0.513\pm 0.16$
ReSample [19] $\bm{0.279}\pm 0.07$ $\textbf{26.36}\pm 2.90$ $0.717\pm 0.11$
DCDP-LDM-DDIM (Ours) $\underline{0.286}\pm 0.08$ $25.97\pm 3.23$ $\underline{0.735}\pm 0.12$
DCDP-LDM-Tweedie (Ours) $0.302\pm 0.09$ $\underline{26.34}\pm 3.34$ $\textbf{0.747}\pm 0.12$
Table 3: Quantitative results of nonlinear deblurring on the ImageNet dataset for both pixel space diffusion models and LDMs. The results colored in blue are for LDMs, while the others are for pixel space diffusion models. The best results are in bold and second best results are underlined for pixel space models and LDMs respectively.
@@ 4.3 Ablation Studies
Finally, we conduct ablation studies demonstrating computational efficiency, convergence, and robustness of the proposed DCDP method.
Figure 7: Comparison of inference times on Motion Deblurring. The experiment setting is the same as that introduced in Section 4.1 and all methods are run using the exact same compute. Here V1 and V2 stands for Approach I and Approach II introduced in Section 3.3 respectively. The inference time for our algorithm is the fastest, while enjoying similar (or better) performance. The runtime is normalized by dividing by the largest value, which corresponds to DPS with a total runtime of 387 seconds.
@@
Computational Efficiency. First, we compare the inference speed (the time it takes to reconstruct a single image) of our method with the baselines and display the results in Figure 7. Clearly, for pixel-based diffusion models, our methods are much more efficient than DPS [14] and MCG [11], as we do not require backpropagation through the score function for data consistency. Furthermore, using the single-step Tweedie’s formula approximation for diffusion purification (DCDP-Tweedie and DCDP-LDM-Tweedie) is much faster than using the reverse ODE (DDIM) which involves multiple reverse diffusion steps (DCDP-DDIM and DCDP-LDM-DDIM). Second, for LDMs, our methods are faster than both ReSample and PSLD. Moreover, enforcing data consistency in the pixel space (Approach II) is considerably more efficient than enforcing data consistency in the latent space (Approach I), as the former does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$.
@@
Figure 8: Convergence of our methods in terms of LPIPS. The orange, green, red and blue curves correspond to the reconstruction process with data fidelity optimization only, DCDP with a constant (large) purification strength, constant (small) purification strength and a linear decaying purificaiton strength, respectively. The x-axis denotes the total number of data fidelity gradient steps. See Figure 20 in Appendix D for convergence behavior measured using PSNR and SSIM.
@@
Empirical Convergence Study. We further investigate the convergence of our algorithm in terms of LPIPS. This study also serves to demonstrate the efficiency of a linear decaying time schedule for $t_{k}$ in diffusion purification. As shown in Figure 8, at the beginning of each diffusion purification and data fidelity optimization cycle, the LPIPS increases after performing diffusion purification due to stochasticity introduced by the forward-backward diffusion process. However, the subsequent iterations successfully decrease LPIPS and result in reconstructions with higher quality compared to pure data fidelity optimization, implying that the purified image is able to serve as a good initialization point. In the same figure, note that the linear decaying schedule greatly improves the performance of our algorithm compared to using a constant time schedule for $t_{k}$. This is expected, as a large constant purification strength will result in serious hallucination while a small constant purification strength is not able to fully alleviate the artifacts in the data fidelity based estimations. We note that similar convergence behavior holds when measuring the reconstruction quality with PSNR and SSIM, as shown in Figure 20 in Appendix D. Lastly, in Figure 9, we illustrate an example reconstruction process for super-resolution with noisy measurements, with more qualitative examples in the Appendix. It is evident that the quality of the reconstructed image gradually increases with each iteration, implying that the diffusion purification is an effective operator for enforcing the image priors embedded in the pre-trained diffusion models.
@@
Metric DPS DCDP-DDIM DCDP-Tweedie DPS DCDP-DDIM DCDP-Tweedie
@@
Noise Robustness. In many practical settings of interest, the measurements $\bm{y}$ can be corrupted with additive noise. Here, we perform an ablation study on the performance of our algorithm in which the measurements have additive Gaussian noise $\mathcal{N}(0,\sigma^{2}_{\bm{y}}\bm{I})$ with varying values of $\sigma_{\bm{y}}$ for super-resolution and motion deblurring. We compare our algorithm to DPS, as their theory states that the approximation of the likelihood gradient becomes tighter as the measurement noise $\sigma_{\bm{y}}$ increases [14]. We conduct experiments on $100$ randomly selected images from the LSUN-Bedroom dataset for super-resolution, motion deblurring with Gaussian measurement noise of standard deviations $0.05,0.075$, and $0.1$. For the hyperparameters, we tune the parameters according to $\sigma_{\bm{y}}=0.1$ and fix them throughout the experiments. The rest of the setup is the same as previously described. We present the results in Table 4.
Note that as the standard deviation increases, the performance of all algorithms degrades. However, our algorithm consistently outperforms DPS in metrics such as PSNR and SSIM, whereas DPS often exhibits higher perceptual quality measured by LPIPS. This difference arises because DPS frequently suffers from hallucinations, introducing artifacts in the image that “sharpen” it. Our algorithm is less susceptible to such artifacts, resulting in “smoother” images with higher signal-to-noise ratios. Details of the hyperparameters and visual results can be found in Appendix C.
Impact of Number of DDIM Steps. Next, we investigate how the number of DDIM steps for diffusion purification affects the reconstruction quality. We focus on solving the nonlinear deblurring with pixel space diffusion models (DCDP-DDIM), varying number of DDIM steps among $\{10,20,30\}$, while fixing all other hyperparameters (see Appendix D for details). As shown in Figure 10(a), increasing the number of DDIM steps degrades PSNR and SSIM—metrics that primarily capture data fidelity—while improving the LPIPS score, which better reflects perceptual quality. This reveals a trade-off between fidelity and perceptual quality controlled by the number of DDIM steps. Based on this observation, we fix the number of DDIM steps to 20 in all DCDP-DDIM and DCDP-LDM-DDIM experiments to strike a balance between the two.
Figure 10: Impact of number of DDIM steps and total number of iterations $K$. Figure (a) illustrates the performance of DCDP-DDIM with different number of DDIM steps for diffusion purification. Figure (b) and (c) illustrate the impact of the total number of iterations $K$ on the performance of DCDP-DDIM and DCDP-Tweedie respectively.
Impact of Total Number of Iterations $K$. Lastly, we investigate how the number of total iterations $K$ affects reconstruction quality. Again, we focus on solving the nonlinear deblurring using pixel space diffusion models (both DCDP-DDIM and DCDP-Tweedie), varying $K$ over $\{5,10,20,30,40\}$ while keeping all other hyperparameters fixed (see Appendix D for details). In each iteration, we use a fixed number of steps $\tau=100$ for data fidelity optimization, resulting in a total of $\{500,1000,2000,3000,4000\}$ data consistency steps. As shown in Figure 10(b) and (c), increasing $K$ initially improves performance across standard metrics, but the gains plateau or even degrade beyond a certain point. In practice, we find that choosing $K$ between 10 and 20 strikes a good balance between reconstruction quality and inference speed.
5 Conclusion
In this study, we introduce a novel diffusion-based method for solving general inverse problems, which decouples the data consistency steps from the reverse diffusion process. Our technique alternates between a reconstruction phase that enforces data consistency and a refinement stage that uses diffusion purification to enforce the diffusion image prior. We demonstrated the effectiveness of our algorithm across a wide range of inverse problems, including deblurring, inpainting, and super-resolution. Notably, by decoupling the two processes, we showed that our algorithm can incorporate a broader range of diffusion methods, such as accelerated samplers, latent diffusion models, and consistency models.
Acknowledgements
QQ, XL, SMK acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, MICDE Catalyst Grant, and a gift grant from KLA. SL and SR acknowledge support from NSF CCF-2212065.
Data Availability Statement
@@ Acknowledgements
Appendix A Discussion on Hyperparameters
@@
Here, we discuss the hyperparameters of our algorithm and some typical choices we make throughout our experiments.
@@
  • Total Number of Iterations. We run the algorithm for a total of $K$ iterations, which refers to reconstructing the image by alternately applying data fidelity optimization and diffusion purification for $K$ rounds. Empirically, we observe a $K$ ranging from 10 to 20 is sufficient for generating high quality reconstructions. We suggest users start with $K=10$ and then gradually decrease it if a faster computation speed is required or increase it if higher reconstruction quality is required.
@@
  • Optimizer for Data Fidelity Optimization. Throughout our experiments, we use momentum-based gradient descent with the momentum hyperparameter set to 0.9. The optimal learning rate ($lr$) for data fidelity optimization varies from task to task. In general, $lr$ should neither be too large, which will not result in images with meaningful structures, nor too small, in which case the convergence speed of the overall algorithm will be too slow. Throughout the experiment we fix the number of gradient steps for each round of the data fidelity optimization. When the measurements are corrupted by noise, one can further perform early stopping to prevent overfitting to the noise.
@@
  • Decaying Time Schedule for Diffusion Purification. The decaying time schedule plays the key role in the success of our proposed algorithm. As described in  Section 3.2, given a data fidelity based reconstruction $\bm{x}_{k}$, the goal of diffusion purification is to generate $\bm{v}_{k}$ that shares the overall structural and semantic information of $\bm{x}_{k}$, such that it is consistent with the measurements while at the same time being free from the artifacts visible in $\bm{x}_{k}$. Unfortunately, the $\bm{v}_{k}$ resulting from diffusion purification is sensitive to the choice of $t_{k}$, i.e., the amount of noise added to $\bm{x}_{k}$. When $t_{k}$ is small, the resulting $\bm{v}_{k}$ will exhibit similar structure to $\bm{x}_{k}$ but with artifacts, as there is not enough noise to submerge these artifacts through diffusion purification. On the other hand, when $t_{k}$ is too large, though $\bm{v}_{k}$ will be free from artifacts, its semantic information and overall structure will significantly deviate from $\bm{x}_{k}$. As a result, the resulting reconstructions will not be measurement-consistent.
@@
    To this end, previous works that utilize diffusion purification [29, 12] depend on fine-tuning an optimal $t^{\star}$. However, this fine-tuning strategy can be infeasible since the optimal $t^{\star}$ varies from task to task and even from image to image. To alleviate the need to identify the optimal $t^{\star}$, we implement a linear decaying schedule $\{t_{k}\}_{k\in[K]}$ for each iteration $k$. In the early iterations when $t_{k}$ is relatively large, the resulting $\bm{v}_{k}$, though inconsistent with the measurements, will be free from artifacts and can serve as a good starting point for the subsequent iteration of data fidelity optimization. As $t_{k}$ keeps decreasing, the algorithm will converge at an equilibrium point which is both consistent with the measurements and free from the artifacts. When the measurements are noiseless, we observe that for a diffusion model trained with 1000 total timesteps, a linear decaying time schedule starting from $t_{1}=400$ and ending at $t_{K}=0$ results in reconstructions with high quality for Super-Resolution, Gaussian Deblurring, and Motion Deblurring. For Box Inpainting, we choose a higher initial timestep $t_{1}=700$ and $t_{1}=500$ for pixel space diffusion models and LDMs respectively to compensate for the severe artifacts in the large box region. When the measurements are corrupted by noise, one may instead end at $t_{K}=\tilde{t}$, where $\tilde{t}>0$, to compensate for the negative effect of the measurement noise. We demonstrate the reconstruction process for various tasks in Figures 11, 12, 13, 14 and 19. Notice that the diffusion purification removes the artifacts from the initial reconstructions obtained via data fidelity optimization, especially in the early iterations, and results in high-quality reconstructions consistent with the measurements as the timestep decreases.
@@
  • DDIM Steps for Diffusion Purification. With the help of accelerated diffusion samplers such as DDIMs [8], we can perform diffusion purification with as few as 20 unconditional sampling steps. As mentioned before, the total number of iterations $K$ ranging from 10-20 is sufficient for generating high quality reconstructions. This amounts to 200 to 400 unconditional reverse sampling steps in total, a significantly lower figure compared to DPS [14], which requires 1000 unconditional reverse sampling steps. Interestingly, we find that increasing the DDIM sampling steps does not necessarily improve the performance of diffusion purification. Therefore, we suggest users use 20 steps for diffusion purification.
  • Pixel Space Optimization and Latent Space Optimization. As discussed in Section 3.4, in the case of LDMs, one can use either Pixel Space Optimization (30) or Latent Space Optimization (27) to enforce the data consistency. Depending on the specific task, one approach may generate $\bm{x}_{k}$ with higher quality than the other. In practice, as shown in Figures 15 and 16, for motion deblurring and Gaussian deblurring, the initial reconstructions generated with Pixel Space Optimization have significantly higher quality compared to the ones generated with Latent Space Optimization. Furthermore, Pixel Space Optimization is much faster than the Latent Space Optimization since the gradient computation of the former approach does not require backpropagation through the deep decoder $\mathcal{D}(\cdot)$. To achieve the best reconstruction quality, we suggest users try both options and choose the one which gives the best performance. On the other hand, for applications that require fast reconstruction speed, we suggest using Pixel Space Optimization.
Figure 15: Comparison between Pixel Space Optimization and Latent Space Optimization for Motion Deblurring. Initial reconstructions obtained via Pixel Space Optimization have much better quality compared to the ones obtained from Latent Space Optimization which in turn result in final reconstructions with higher quality.